A monopolist sells to two consumer groups, students and non-students.

   Demand for students: Q = 500 - 1/2P

      Demand for non-students: Q = 750 - 2P

      MC = 20

Find the profit-maximizing price/quantity combination in each market if the groups can be separated.
Answer #1

Market for students

Profit = revenue - cost

πstudents = (p - c)q

= (p - 20)(500 - 0.5p)

= 500p - 1000 - 0.5p^2 + 10p

= 510p - 0.5p^2 - 1000

Profit is maximum when π'(p) = 0

510 - p = 0

p = $510

q = 500 - 510/2 = 245 units

Hence, profit-maximizing price/quantity combination in market for students is P = $510 and Q = 245 units

Market for non students

Profit = revenue - cost

πnon-students = (p - c)q

= (p - 20)(750 - 2p)

= 750p - 1500 - 2p^2 + 40p

= 790p - 2p^2 - 1500

Profit is maximum when π'(p) = 0

790 - 4p = 0

p = $197.50

q = 750 - 197.50*2 = 355 units

Hence, profit-maximizing price/quantity combination in market for non students is P = $197.50 and Q = 355 units